OpenStudy (anonymous):

lim as x approaches 25 for.. (sqrt(x)-5)/(x-25) = ?

hartnn (hartnn):

hint : \(\large x-25 = (\sqrt{x})^2-5^2\) now use the fact that \(a^2-b^2=(a+b)(a-b)\) to factor that.

OpenStudy (anonymous):

Thankies

hartnn (hartnn):

welcome ^_^ if you want to verify your answer, i'll be there.

OpenStudy (anonymous):

Alright, :b I got 1/10. Is that correct?

hartnn (hartnn):

absolutely :)